What comes next- 4,5,9,14,23,37?
Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

Another hurricane.

Answer by Theo(13342)

the sequence is as follows:

4 + 5 = 9
5 + 9 = 14
9 + 14 = 23
14 + 23 = 37
23 + 37 = 60

the next number in the sequence should be 60.

the pattern looks to be like this geometrically:



4 + 5 = 9
    5 + 9 = 14
        9 + 14 = 23
            14 + 23 = 37
                 23 + 37 = 60


you next number should be 60 if i figured this out correctly.